a) O X(t) = 10 + 3t2, y(t) = 4, t€ (-7,-3] b) O x(t) = 12, y(t) = 13 + 3t +4, + € [49,9] c) O x(t) = t, y(t) = 40 + 372 +4, tPLEASE EXPLAIN
Answer #1

Given that fay=x6 +30²74 : Now wehere to find a parametriction x= x(t), y = y(t) for f(x)= x6 + 3x44 from 6-74-10) to (-313)